OpenStudy (anonymous):

determine the vertex (both coordinates). (-x^2/3)-(x/2)-(1/5)

OpenStudy (amistre64):

1/2 ---- = -3/4 is your x coord i believe -2/3

OpenStudy (amistre64):

-0.0125 = -1/8 would be the y coord.... if i see it right

OpenStudy (amistre64):

(-1/3)x^2 -(1/2)x -(1/5) -2/3x -1/2 x = -1/2//2/3 = -3/4 ....yep

OpenStudy (amistre64):

2.5 4.25 8.125 16.0625

OpenStudy (amistre64):

-5/16 maybe

OpenStudy (amistre64):

125/1000 lol

OpenStudy (anonymous):

\[\frac{-b}{2a}=\frac{\frac{1}{2}}{2\times -\frac{1}{3}}=-\frac{3}{4}\]\]

OpenStudy (amistre64):

gonna have to ask a brit

OpenStudy (amistre64):

(-.75,-.0125) ;)

OpenStudy (anonymous):

i got -.95

OpenStudy (amistre64):

(-((-.75)^2)) / 3) - ((-.75) / 2) - (1 / 5) = -0.0125 hot of the google

OpenStudy (anonymous):

\[-2\times (-\frac{3}{4})^2-\frac{1}{2}\times -\frac{3}{4}-\frac{1}{5}\]

OpenStudy (amistre64):

-2?

OpenStudy (anonymous):

you are right. i was multiplying the first term by -2 because i am not paying attention.

OpenStudy (amistre64):

it helps to sit facing a wall; that way when the butterflys float by you dont get distracted lol

OpenStudy (anonymous):

just came to see if you were still here. you are more obsessed than i am!

OpenStudy (anonymous):

thanks guys

OpenStudy (anonymous):

im stiil trying to figure out you plug in -3/4 for x right?

OpenStudy (amistre64):

yes, x = -3/4 = -.75

OpenStudy (anonymous):

and how is y?

OpenStudy (amistre64):

you simply use the value for x in the equation to determine what 'y' equals.... no easy way around htat